2. Change the total fixed manufacturing overhead cost for the Milling Department in Data area back to $390,000, keeping all oRestore the total number of machine-hours in the Assembly Department to 3,000 machine-hours. And keep the job data the same as it was in Requirement 2. What is the selling price for Job 408 if the total number of direct labor-hours in the Assembly Department decreases from 80,000 direct labor-hours to 40,000 direct labor-hours?
Answer #1

Selling Price of Job 408 = $5451.25

Markup on job cost 75%
Department
Milling Assembly
Machine hours 60000 3000
Direct Labor hours 8000 40000
Total Fixed Manufacturing Overhead $      390,000 $         500,000
Variable manufacturing overhead per machine hour $            2.00
Variable manufacturing overhead per direct labor hour $                3.75
Cost Summary for Job 408 Department
Milling Assembly
Machine Hours 70 3
Direct Labor hours 5 40
Direct Material $              830 $                 320
Direct Labor Cost $                80 $                 640
Milling Assembly
Total Fixed Manufacturing Overhead $      390,000 $         500,000
Variable manufacturing overhead per machine hour $            2.00 $                3.75
Total Machine hours or direct labor hours 60000 40000
Total Variable Manufacturing Overhead $      120,000 $         150,000
Total Manufacturing Overhead $      510,000 $         650,000
Milling Assembly
Total Manufacturing Overhead $      510,000 $         650,000
Total Machine hours or direct labor hours 60000 40000
Predetermined Overhead rate $            8.50 $              16.25
Milling Assembly
Predetermined Overhead rate $            8.50 $              16.25
Machine hours or direct labor hours for the job 70 40
Manufacturing overhead applied $              595 $                 650
Milling Assembly Total
Direct Materials $              830 $                 320 $            1,150
Direct Labor cost $                80 $                 640 $                720
Manufacturing overhead applied $              595 $                 650 $            1,245
Total Cost of Job 408 $            3,115
Total Cost of Job 408 $       3,115.00
Markup $       2,336.25
Selling Price of Job 408 $      5,451.25

Formula Sheet

Milling Assembly
Total Fixed Manufacturing Overhead =B10 =C10
Variable manufacturing overhead per machine hour =B11 =C12
Total Machine hours or direct labor hours =B8 =C9
Total Variable Manufacturing Overhead =B26*B27 =C26*C27
Total Manufacturing Overhead =B25+B28 =C25+C28
Milling Assembly
Total Manufacturing Overhead =B29 =C29
Total Machine hours or direct labor hours =B27 =C27
Predetermined Overhead rate =B33/B34 =C33/C34
Milling Assembly
Predetermined Overhead rate =B35 =C35
Machine hours or direct labor hours for the job =B16 =C17
Manufacturing overhead applied =B39*B40 =C39*C40
Milling Assembly Total
Direct Materials =B18 =C18 =B45+C45
Direct Labor cost =B19 =C19 =B46+C46
Manufacturing overhead applied =B41 =C41 =B47+C47
Total Cost of Job 408 =D45+D46+D47
Total Cost of Job 408 =D48
Markup =D51*B4
Selling Price of Job 408 =D51+D52

If you have any query, kindly comment with your query and please mark thumbs up.